If pressure is expressed in atmospheres and volume is expressed in liters, PΔV has units of L atm (liters × atmospheres). Remember that 1 atm = 101,325 Pa, and that 1 L = 1 dm3. Use this information to determine the number of joules corresponding to 2 L atm.


1
Expert's answer
2022-05-05T12:51:03-0400

"J=Pa (m^3)=1000 Pa(l)=1000\/101325 atm(l)=0.0099 atm(l)"

2 L atm=2/0.0099=202 J